FOURIER TRANSFORM INFRARED EMISSION SPECTRA OF CaH AND SrH

A. SHAYESTEH, I. GORDON, D. R. T. APPADOO, \underline{K. A. WALKER}, and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada.

The vibration-rotation spectra of gaseous CaH and SrH have been recorded in emission using a furnace-discharge source. Ca and Sr vapor were produced at 780^{o}C and 850^{o}C, respectively, and reacted with a few torr of hydrogen gas under a 3 kV / 333 mA discharge to produce the metal hydrides. The fundamental and first three hot bands, v=1-0 to v=4-3, in the ^{2}Σ^{+} ground electronic state of each molecule were observed. The analysis of the data is in progress and will be presented.